OpenStudy (anonymous):

Hi have a 4 step question 1. To open these doors, you must speak three functions in standard form. One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions. 2. Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. You may use the f(x) you created in question 1 as an example. 3. Find the solutions of g(x). Show each step. 4. Justify if completing the square is a good method for solving when the Discriminant is negative

OpenStudy (anonymous):

please help

hero (hero):

Start with f(x) = (x + a)(x + b) where a and b are two integers such as 3 and 4 g(x) = (x + a)(x + b) where a and b are two irrational numbers such as sqrt(2) or pi h(x) = (x + a)(x + b) where a and b are two complex numbers such as 3 + 4i , 3 - 4i Then multiply each one out to put them in standard form